OpenStudy (anonymous):

Jeff and Lucy have been asked to wash their mom’s minivan. It takes Jeff 2 hours to wash the van by himself, and it takes Lucy 1.5 hours to wash the van by herself. How long will it take Jeff and Lucy to wash the van if they work together? A:What equation is used to solve this problem? What does each variable represent? B:Identify the given information. C:Enter the given information into the equation. D:Solve the equation. E:State the solution.

OpenStudy (saranya):

Jeff takes 2 hours to wash the car. Thus, Jeff's 1-hour work =1/2 the car-wash Lucy takes 3/2 hours to wash the car. Thus, Lucy's 1-hour work =1/(3/2)=2/3 the car-wash Now, (Jeff+Lucy)'s 1-hour work =(1/2)+(2/3)=7/6 the wash Hence, time taken by Jeff+Lucy to finish the car-wash =1/(7/6) =6/7 hour I hope this helps yu ! :)